# FairShare: Auditable Geographic Fairness for Multi-Operator LEO Spectrum Sharing

### v1

@@ VI Results
Figure 3: SNR distribution by geographic category. Urban users experience a systematic 5–8 dB advantage due to favorable elevation geometry.
@@ VI-A Geographic Allocation
Before examining allocation outcomes, Fig. 3 confirms the structural SNR gap underlying our fairness analysis: urban users exhibit a median SNR approximately 6 dB higher than rural users. This physical-layer disparity directly translates to allocation bias under channel-aware policies. Table I presents the allocation performance over $n=50$ runs.
Key Findings: Priority scheduling yields a $1.65\times$ disparity. More critically, it results in a service outage probability of $73.3\%$ for rural users (since $\rho_{\text{rural}}=26.7\%$), compared to only $56\%$ outage for urban users. FairShare (with 35% rural quota) reduces the rural outage probability to $66.7\%$, effectively bridging the service availability gap.
TABLE I: Allocation rates and fairness ($W=300$ MHz, $n=50$). $\Delta_{\text{geo}}=\rho_{\text{urban}}/\rho_{\text{rural}}$; =1=1: perfect fairness, >1>1: urban bias, <1<1: rural bias.
@@
Equal Static 77.9 $\pm$ 0.4% 78.1 $\pm$ 0.7% 1.00 $\pm$ 0.01
SNR Priority 44.0 $\pm$ 0.0% 26.7 $\pm$ 0.0% 1.65 $\pm$ 0.00
Demand Prop. 40.8 $\pm$ 0.0% 29.1 $\pm$ 0.0% 1.40 $\pm$ 0.00
FairShare 24.0 $\pm$ 0.0% 33.3 $\pm$ 0.0% 0.72 $\pm$ 0.00
Bandwidth Sensitivity: Table II(a) reveals a counter-intuitive trend: disparity under Priority scheduling increases with bandwidth (from $1.00\times$ at 50 MHz to $1.65\times$ at 300 MHz). In highly constrained regimes (50 MHz), the system is saturation-limited, treating all users equally poorly. As capacity grows, the scheduler exercises its preference for high-SNR (urban) users, revealing the latent structural bias.
@@
Quota Tunability: Table II(b) demonstrates FairShare’s flexibility as a policy tool. By adjusting the rural quota from 25% to 40%, $\Delta_{\text{geo}}$ shifts from urban-biased ($1.11\times$) to strongly rural-favoring ($0.50\times$). The default 35% quota yields $\Delta_{\text{geo}}=0.72\times$, representing affirmative action for underserved rural users. This tunability allows regulators to set precise fairness targets based on policy objectives.
Robustness to Channel Assumptions: Table III evaluates an extreme scenario with inverted SNR correlation (rural users have higher SNR). Equal Static remains neutral ($\Delta_{\text{geo}}\approx 1$) but inefficient. We also test a Demand-Weighted variant (Demand Proportional with 3:1 urban-to-rural demand weighting) modeling aggressive commercial bias. Key insights: (1) SNR Priority shifts to favor rural users ($0.76\times$); (2) Demand Proportional still favors urban users ($1.22\times$); (3) Demand-Weighted shows catastrophic urban bias ($3.38\times$ and $2.11\times$), proving commercial incentives can override SNR advantages; (4) FairShare maintains consistent fairness ($0.72\times$), confirming quota robustness against both physical and policy-induced biases.
@@
TABLE II: Sensitivity analysis. All disparity values are urban-to-rural ratios $\Delta_{\text{geo}}=\rho_{\text{urban}}/\rho_{\text{rural}}$ ($\times$).
(a) Impact of Bandwidth on Geographic Disparity
Bandwidth Equal Priority Demand FairShare
50 MHz 1.00 1.00 1.00 0.72
100 MHz 1.00 1.00 1.00 0.72
@@
(b) Impact of Rural Quota (FairShare Only)
Quota $\boldsymbol{\Delta_{\text{geo}}}$ Policy Interpretation
@@
TABLE III: Policy behavior under inverted SNR correlation: urban vs. rural SNR advantage.
Policy Standard Inverted
(Urban SNR+) (Rural SNR+)
[…]
FairShare $\times 0.72$$\times$
Demand-Weighted: Demand Proportional with 3:1 urban-to-rural demand weighting, modeling extreme commercial bias.
@@ VII Discussion
VII-A The Illusion of Aggregate Fairness
Conventional fairness metrics, while suitable for terrestrial networks, obscure geographic disparities of NTNs. Although all evaluated policies achieve Jain’s index $J>0.95$, SNR-priority scheduling yields a $1.65\times$ urban-rural disparity, which means that urban users are 65% more likely to receive service. This bias arises because aggregate metrics are inflated by allocations to the dominant urban population (50%). Our results challenge the neutrality of channel-aware scheduling in NTNs, revealing it as an inadvertent mechanism for geographic discrimination. This underscores the necessity of geographic-specific metrics, such as $\rho_{\ell}$ and $\Delta_{\text{geo}}$, for meaningful fairness auditing.
VII-B Policy Design Over Infrastructure Scaling
Figure 4: Efficiency-fairness tradeoff: FairShare achieves near-optimal throughput while ensuring $\Delta_{\text{geo}}<1$.
The bandwidth sensitivity analysis (Table II(a)) confirms that geographic unfairness is policy-inherent, not scarcity-driven. As bandwidth increases from 50 to 300 MHz, the disparity under priority scheduling grows from $1.0\times$ to $1.65\times$, exposing structural bias that relaxed constraints unmask.
Scalability Analysis: As shown in Table IV-top, varying the number of operators from 1 to 10 alters priority-based disparity by less than $\pm 10\%$. Increasing satellite density per operator from 20 to 100 provides only a marginal $\sim$0.5 dB SNR gain for rural users (Table IV-bottom), dwarfed by the structural 6–8 dB penalty from elevation geometry (Fig. 3). These results demonstrate that unfairness is robust to network scaling; it is intrinsic to allocation policy, not an artifact of resource scarcity.
TABLE IV: Scalability analysis: Geographic disparity ($\Delta_{\text{geo}}$) across network scales.
Configuration $\boldsymbol{\Delta_{\text{geo}}}$
# Ops Spec./Op. Priority FairShare
1 1000 MHz 1.54 0.71
3 333 MHz 1.65 0.80
5 200 MHz 1.47 0.60
10 100 MHz 1.40 0.60
Sats/Op. (3 ops) $\boldsymbol{\Delta_{\text{geo}}}$ (Priority)
20 (Sparse) 1.47
50 (Medium) 1.47
100 (Dense) 1.46
[…]
@@ VII-C FairShare as a Regulatory Tool
FairShare establishes the Pareto frontier for fair resource allocation. As Proposition 1 states, for any geographic quotas $\{q_{\ell}\}$, FairShare maximizes the sum-rate while guaranteeing fairness. Our results confirm this: FairShare achieves affirmative disparity ($\Delta_{\text{geo}}=0.72\times$) while maintaining high efficiency and reducing runtime by 3.3%. The partition-then-optimize design offers regulators a practical tool with three advantages: (1) transparency (quotas are interpretable and auditable), (2) efficiency (no throughput penalty), and (3) practicality (reduced runtime enables real-time operation). We propose mandating the reporting of $\rho_{\ell}$ and $\Delta_{\text{geo}}$, establishing minimum access guarantees and transparent policies, and adopting tunable quota frameworks such as FairShare to set enforceable fairness targets (e.g., $\Delta_{\text{geo}}\leq 1.2\times$).
VII-D Boundaries and Future Work
Our research employs controlled assumptions to isolate structural geometric bias, defining its scope and future directions. The NYC-centric, full-buffer model provides a conservative baseline; in continental deployments, elevation disparities would be more severe. While orbital dynamics alter instantaneous visibility, the long-term statistical bias remains.
FairShare assumes a trusted central coordinator, aligning with regulatory intervention. Future work needs to design incentive-compatible mechanisms for competitive operator markets and investigate dynamic quota adaptation. Extensions to uplink scenarios, interference modeling, and validation with real LEO measurements are future research directions.
From Resource to Experience: FairShare ensures fair access to spectrum rather than throughput optimization. In congested NTNs, resource access is the primary bottleneck; without allocation, users experience complete outage. By guaranteeing $\Delta_{\text{geo}}\approx 1$, FairShare equalizes the fundamental probability of service access, establishing a first-order baseline for equitable experience upon which higher-layer enhancements can be devised.
@@ VIII Conclusion
This research has established that conventional spectrum allocation policies are fundamentally misaligned with geographic equity in multi-operator NTNs. Through 3GPP-compliant simulations, we have demonstrated that SNR-priority scheduling creates a $1.65\times$ urban-rural access disparity, a bias that is policy-inherent and exacerbated, not alleviated, by increasing the bandwidth or infrastructure scale. We have introduced FairShare, a lightweight quota-based framework that achieves affirmative fairness ($\Delta_{\text{geo}}=0.72\times$) while maintaining efficiency and reducing computational overhead. The core insight of our analysis is that explicit geographic partitioning is both necessary and sufficient for equity. By providing the diagnostic metric $\Delta_{\text{geo}}$ and an enforceable quota mechanism, FairShare transforms a technical scheduling problem into actionable policy, equipping regulators with the tools needed to mandate equitable spectrum access and close the digital divide.

### v2

@@ VI-A Geographic Allocation
The structural SNR gap (urban users exhibit $\sim$6 dB higher median SNR than rural users due to favorable elevation geometry) directly translates to allocation bias under channel-aware policies. Table II presents the allocation performance for the Starlink Shell 1 constellation, averaged over $20\times 50=1{,}000$ samples.
Key Findings: Priority scheduling yields a $1.84\times$ mean disparity with substantial temporal fluctuations (std =0.93=0.93), peaking at $3.9\times$ during low-interference orbital configurations. It results in a mean service outage probability of $74.3\%$ for rural users ($\rho_{\text{rural}}=25.7\%$), compared to $60.7\%$ for urban users. FairShare (with 35% rural quota) reduces the rural outage probability to $59.0\%$, effectively bridging the service availability gap with zero temporal variance.
TABLE II: Allocation rates and fairness (Starlink Shell 1, $W=300$ MHz, $20\times 50$ samples, with interference). $\Delta_{\text{geo}}=\rho_{\text{urban}}/\rho_{\text{rural}}$; std reflects temporal variability.
@@
Equal Static 35.2 $\pm$ 0.2% 35.2 $\pm$ 0.3% 1.01 $\pm$ 0.01
SNR Priority 39.3 $\pm$ 3.9% 25.7 $\pm$ 8.3% 1.84 $\pm$ 0.93
Demand Prop. 38.7 $\pm$ 0.2% 29.8 $\pm$ 0.3% 1.31 $\pm$ 0.02
FairShare 28.0 $\pm$ 0.0% 41.0 $\pm$ 0.0% 0.68 $\pm$ 0.00
Figure 3: Geographic disparity ($\Delta_{\text{geo}}$) over 20 orbital snapshots (Starlink Shell 1). SNR Priority exhibits large temporal fluctuations driven by interference geometry changes, while FairShare remains perfectly flat at $0.68\times$.
Temporal Variability: A striking finding, visualized in Fig. 3, is the large temporal variability of SNR Priority ($\text{std}=0.93$) compared to FairShare ($\text{std}=0.00$). As satellites orbit, the beam overlap pattern changes, causing co-channel interference to oscillate with an approximately 5-minute period. The Pearson correlation between average SINR and $\Delta_{\text{geo}}$ for Priority scheduling is $r=0.69$: when interference is low (SINR $\approx$ 28 dB), the urban–rural quality gap is amplified and disparity peaks at $3.9\times$; when interference is high (SINR $\approx$ 10 dB), it partially equalizes users, reducing disparity to $\sim$$1\times$. FairShare’s quota mechanism renders it completely invariant to these dynamics the allocation counts per region are deterministic, yielding $\Delta_{\text{geo}}=0.68\times$ at every snapshot. This demonstrates that orbital dynamics expose priority-based policies as not only unfair on average but temporally unstable, while FairShare provides guaranteed fairness.
Cross-Constellation Validation: Table III demonstrates that the structural bias is not an artifact of a specific constellation geometry. All three constellations spanning altitudes from 550 to 1,200 km and sizes from 648 to 1,584 satellites exhibit significant urban bias under Priority scheduling, with $\Delta_{\text{geo}}$ ranging from $1.60\times$ to $1.84\times$. FairShare achieves an identical $\Delta_{\text{geo}}=0.68\times$ across all three, confirming that its quota mechanism is robust to constellation design.
TABLE III: Cross-constellation comparison of geographic disparity.
$\boldsymbol{\Delta_{\text{geo}}}$
Constellation Alt. Sats Priority FairShare
Starlink Shell 1 550 km 1,584 1.84 $\pm$ 0.93 0.68 $\pm$ 0.00
OneWeb Phase 1 1,200 km 648 1.77 $\pm$ 0.59 0.68 $\pm$ 0.00
Kuiper Shell 1 630 km 1,156 1.60 $\pm$ 0.49 0.68 $\pm$ 0.00
Interference Impact: Table IV compares system behavior with and without inter-beam co-channel interference. The 4-color frequency reuse and hexagonal beam layout reduce SINR by approximately 15 25 dB relative to interference-free SNR, placing users in a realistic operating regime. Notably, interference moderately reduces the mean disparity for Priority scheduling (from $2.55\times$ to $1.84\times$) because co-channel interference partially equalizes channel conditions across regions. FairShare remains invariant ($0.68\times$) regardless of interference, confirming its robustness.
TABLE IV: Impact of inter-beam co-channel interference (Starlink Shell 1). Average SINR is computed over allocated users only.
Avg. SINR (dB) $\boldsymbol{\Delta_{\text{geo}}}$
Policy No Intf. With Intf. No Intf. With Intf.
Equal Static 42.3 20.8 1.00 1.01
SNR Priority 47.2 32.7 2.55 1.84
Demand Prop. 42.9 22.0 1.36 1.31
FairShare 46.6 32.1 0.68 0.68
Bandwidth Sensitivity: Table V(a) reveals that disparity under Priority scheduling increases with bandwidth. At 200 MHz, the structural bias emerges ($1.29\times$); at 300 MHz it reaches $1.65\times$. Below 200 MHz, the system is saturation-limited and all policies yield $\Delta_{\text{geo}}=1.00$. FairShare maintains $0.72\times$ regardless of bandwidth.
@@
Quota Tunability: Table V(b) demonstrates FairShare’s flexibility: adjusting the rural quota from 25% to 40% shifts $\Delta_{\text{geo}}$ from urban-biased ($1.11\times$) to strongly rural-favoring ($0.50\times$), enabling regulators to set precise fairness targets.
@@
TABLE V: Sensitivity analysis (single-snapshot, no-interference baseline). $\Delta_{\text{geo}}=\rho_{\text{urban}}/\rho_{\text{rural}}$ ($\times$); consistent with multi-snapshot results (Table II).
(a) Bandwidth Impact
BW Equal Priority Demand FairShare
@@
(b) Rural Quota Tunability
Quota $\boldsymbol{\Delta_{\text{geo}}}$ Interpretation
@@
Achieves $\Delta_{\text{geo}}=0.68$ under multi-snapshot with interference (Table II).
@@ VII Discussion
VII-A Aggregate Metrics Mask Geographic Bias
Although all policies achieve Jain’s index $J>0.95$, SNR-priority scheduling yields $1.84\times$ mean urban-rural disparity up to $3.9\times$ during specific orbital configurations because aggregate metrics are inflated by the dominant urban population (50%). The interference-modulated oscillation in $\Delta_{\text{geo}}$ (Fig. 3) further compounds this: a regulator auditing at a single instant may observe values from $0.9\times$ to $3.9\times$. FairShare’s quota mechanism eliminates both problems, providing $\Delta_{\text{geo}}=0.68\times$ with zero variance at every snapshot. These findings underscore that geographic-specific metrics ($\rho_{\ell}$, $\Delta_{\text{geo}}$) are essential for meaningful NTN fairness auditing.
VII-B Unfairness Is Policy Inherent
The cross-constellation results (Table III) demonstrate that unfairness persists despite fundamentally different orbital parameters, and the bandwidth analysis (Table V(a)) shows disparity grows with capacity proving the bias is structural, not scarcity-driven. As Proposition 1 establishes, FairShare achieves the Pareto-optimal sum-rate within any quota constraint. Deploying more satellites or spectrum alone will not close the digital divide; regulatory interventions must prioritize fair allocation policies.
@@ VII-C FairShare as a Regulatory Tool
The partition-then-optimize design offers regulators transparency (auditable quotas), efficiency (no throughput penalty, 3.3% runtime reduction), and practicality (real-time operation). FairShare is architecturally compatible with existing spectrum governance: in the FCC’s SAS/CBRS framework, a centralized coordinator already enforces tiered access policies. FairShare requires only a geographic classification layer and configurable quota parameters $\{q_{\ell}\}$ no modifications to operators’ PHY or MAC layers. The $\Delta_{\text{geo}}$ metric provides a standardized, time-invariant audit trail aligned with NTIA’s broadband equity mandates [6]. As the FCC extends shared-spectrum frameworks to NGSO services [10], FairShare offers a ready-to-deploy fairness module complementing existing interference management. We propose mandating the reporting of $\rho_{\ell}$ and $\Delta_{\text{geo}}$ and adopting tunable quota frameworks to set enforceable targets (e.g., $\Delta_{\text{geo}}\leq 1.2\times$).
VII-D Scope and Future Work
The Keplerian propagator suffices for our 10-minute window; higher-order perturbations matter only for longer studies. The full-buffer traffic and NYC-centric models provide conservative baselines; extension to bursty traffic, continental coverage, and proportional-fair intra-region scheduling are natural next steps. FairShare assumes a trusted coordinator; designing incentive-compatible mechanisms for competitive markets and validating with real LEO measurements remain open.
@@ VIII Conclusion
Through 3GPP-compliant simulations with Keplerian orbital dynamics, inter-beam interference, and three constellation geometries, we have shown that SNR-priority scheduling creates a $1.84\times$ mean urban-rural disparity (peak $3.9\times$) a bias that is policy-inherent, persists across constellations, and worsens with increasing bandwidth. FairShare, a lightweight quota-based framework, achieves $\Delta_{\text{geo}}=0.68\times$ with zero variance across all conditions while reducing runtime by 3.3%. Its complete invariance to physical-layer dynamics makes it uniquely suited for regulatory enforcement, transforming a scheduling problem into actionable spectrum policy for closing the digital divide.